OpenStudy (haleyelizabeth2017):

Elimination in Alg 2 :P

OpenStudy (haleyelizabeth2017):

-3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9 options: a. (5, -2, 7) b. (-5, 2, 7) c. (5, 2, -7) d. no solution

OpenStudy (gorv):

yeah type up

OpenStudy (haleyelizabeth2017):

Do you know how to do that?

OpenStudy (gorv):

no solution

OpenStudy (haleyelizabeth2017):

thank you so much :)

OpenStudy (gorv):

solve in set of 2

OpenStudy (haleyelizabeth2017):

Okay thank you so much!

OpenStudy (gorv):

wlcm:)